### v2

@@ Appendix A Hyperparameters
Input Size $28\times 28$
@@
FC Layer 1: Input Features $16\times 16\times 16$
@@
Appendix B Results on MNIST
Here we present our results on the MNIST (Deng, 2012) dataset. For the clusterability of the model with $k$ bipartite clusters, see Figure 9, and for the class-wise accuracies for each label with clusters turned ON and OFF, see Figure 10.
@@ Appendix B Results on MNIST
Appendix C Clustered Layer Visualization
Figure 12 shows the visualization of the fully connected layer trained on CIFAR10 of the network with bipartite clusters.
@@ Appendix C Clustered Layer Visualization
Appendix D Interpretability of Modular Transformer for Modular Arithmetic
In Figure 13, we show the $L2$ norms of the embeddings with the discrete Fourier transformation are very similar to the observations made in Nanda et al. (2023), indicating that the circuits in the modular model learn the same algorithm to perform modular addition.
However, interpreting such a model is much easier with modularity, because modular MLPs have a much simpler representation space, and modular Attention matrices imply a $n$ times simpler circuit search space as compared to a non-modular model in the presence of $n$ completely non-intersecting clusters with a clusterability value of $1$, as we observe in LABEL:sec:modular-modular-addition.
@@ Appendix D Interpretability of Modular Transformer for Modular Arithmetic
Appendix E Type-1 and Type-2 Intervention Performance
Figure 14 shows Type $1$ and Type $2$ interventions illustrating the dependence of samples on individual or pairs of modules. The figure highlights a significant difference in sample dependence between modular and non-modular models, particularly in Type $1$ interventions. Approximately $30\%$ of samples in the non-modular model rely on all $4$ modules for correct predictions, compared to only about $1\%$ in the modular model.
Figure 14: Type $1$ and Type $2$ interventions illustrating the dependence of samples on individual or pairs of modules. The figure highlights a significant difference in sample dependence between modular and non-modular models, particularly in Type $1$ interventions. Approximately $30\%$ of samples in the non-modular model rely on all $4$ modules for correct predictions, compared to only about $1\%$ in the modular model. Similarly, in Type $2$ interventions, there is a notable $10\%$ difference in module dependence on all $4$ samples between the two models.
Appendix F Explanation of LLM Intervention Values
@@
Appendix G Clusterability for Pythia Models for Different Number of Clusters
In Table 2, we share the clusterability for pythia-70m on different values of $k$ (number of clusters).
@@ Appendix G Clusterability for Pythia Models for Different Number of Clusters
Appendix H Details on Other Metrics
H.1 Community Structure

### v3

@@ Appendix B Hyperparameters
Input Size $\times 28$ 28
@@
FC Layer 1: Input Features $\times 16\times 16$ 16
@@
Appendix C Results on MNIST
Here we present our results on the MNIST (Deng 2012) dataset. For the clusterability of the model with $k$ bipartite clusters, see Fig. 10, and for the class-wise accuracies for each label with clusters turned ON and OFF, see Fig. 11.
@@ Appendix C Results on MNIST
Figure 13: Combined clusterability of the MLP input of all layers in GPT2-small and the cross-entropy training loss on a subset of the Wikipedia dataset during modularity training.
Figure 14: Intervention results on LLMs on Chemistry (Modular) for two levels of semantic clustering.
Figure 15: Intervention results on LLMs on Chemistry (Non-modular) for two levels of semantic clustering.
Appendix D Clustered Layer Visualization
Fig. 16 shows the visualization of the fully connected layer trained on CIFAR10 of the network with bipartite clusters.
@@ Appendix D Clustered Layer Visualization
Appendix E Interpretability of Modular Transformer for Modular Arithmetic
In Fig. 17, we show the $L2$ norms of the embeddings with the discrete Fourier transformation are very similar to the observations made in Nanda et al. (2023), indicating that the circuits in the modular model learn the same algorithm to perform modular addition.
However, interpreting such a model is much easier with modularity, because modular MLPs have a much simpler representation space, and modular Attention matrices imply a $n$ times simpler circuit search space as compared to a non-modular model in the presence of $n$ completely non-intersecting clusters with a clusterability value of 111, as we observe in LABEL:sec:modular-modular-addition.
@@ Appendix E Interpretability of Modular Transformer for Modular Arithmetic
Appendix F Type-1 and Type-2 Intervention Performance
[…]
Appendix G Explanation of LLM Intervention Values
@@
Appendix H Clusterability for Pythia Models for Different Number of Clusters
In Fig. 2, we share the clusterability for pythia-70m on different values of $k$ (number of clusters).
@@ Appendix H Clusterability for Pythia Models for Different Number of Clusters
Appendix I Chemistry on Gemma
Figure 19: Intervention results on a Chemistry dataset with a clustered Gemma-1B model with 444 clusters.
Appendix J Details on Other Metrics
J.1 Community Structure